OpenStudy (anonymous):

surface area of cylinders height 10 and radius of 10 how do i work this out

OpenStudy (anonymous):

2 π r2 + 2 π r h

OpenStudy (anonymous):

\[surface area of cylinder is =2pir ^{2}+2pirh\] so = 2*pi*100+2*pi*100

OpenStudy (anonymous):

It is a closed cylinder or open cylinder????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what does yes mean ??? open cylinder like a pipe or closed cylinder like a pepsi can??

OpenStudy (anonymous):

closed cylinder see i know to use 2 π r2 + 2 π r h and when i do i get but it give me answer and that is not one of them

OpenStudy (anonymous):

What r your options??

OpenStudy (anonymous):

340 ,150,130,176,384,360,112,720

OpenStudy (anonymous):

Any unit for radius and height like m, cm etc.???

OpenStudy (anonymous):

well see i got it now they were doing it like \[\pi*5^2=25\pi then 2\pi5=10\pi 10\pi*10=100\pi +\pi*5^2=25\pi at the end you add 25\pi+25\pi+100\pi\] and gett 150pi

OpenStudy (anonymous):

From where has the 5 come??? Y said radius was 10 and height was 10